OpenStudy (anonymous):

All of the following are equivalent, except _____. A.) 8x + 3x B.) (8 + 3)x C.) 11x D.) 11x^2

OpenStudy (anonymous):

ignoring gggurumanjunath, what do you think the answer is @xo_kansasprincess_xo

OpenStudy (anonymous):

I'm not sure.

OpenStudy (anonymous):

well, you have 3 answers that are equivalent, and 1 that is not

OpenStudy (anonymous):

do you know the answer of 8x + 3x?

OpenStudy (anonymous):

I think it's 11x^2

OpenStudy (anonymous):

close... but.. x^2 = x*x 2x = x+x

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

in other words, when you multiply a number by itself, you are squaring it when you add a number to itself, its essentially like multiplying by 2

OpenStudy (anonymous):

alright i'll explain geometrically

OpenStudy (anonymous):

the area of a rectangle is found by multiplying the width by the height, yes?

OpenStudy (anonymous):

like